OpenStudy (anonymous):

The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2) y = 1 over 2x + 3 over 2 y = negative 1 over 2x + 3 over 2 y = 1 over 2x − 3 over 2 y = − 1 over 2x − 3 over 2

OpenStudy (shamim):

given y=-2x-5 so slope m1=-2

OpenStudy (shamim):

if slope of a perpendicular line is m2

OpenStudy (anonymous):

it has to be perpendicular so the slope would be \[\frac{ 1 }{ 2}\], but how do i make it so it passes through point (-1, -2)?

OpenStudy (shamim):

we can write m1*m2=-1 -2*m2=-1 m2=1/2

OpenStudy (shamim):

let the ewuation of perpendicular line is y=m2*x+c

OpenStudy (shamim):

need to find out c=?

OpenStudy (shamim):

put x=-1 y=-2

OpenStudy (shamim):

tell me c=?

OpenStudy (abmon98):

-2=-1/2+c c=-4+1 c=-3

OpenStudy (shamim):

also put m2=1/2

OpenStudy (anonymous):

so it would be y - (-2) = 1/2 (x + 1)?

OpenStudy (abmon98):

yes thats right

OpenStudy (shamim):

no

OpenStudy (shamim):

slope intetcept form is y=mx+c

OpenStudy (abmon98):

y-y1=m(x-x1)

OpenStudy (shamim):

-2=(1/2)*(-1)+c

OpenStudy (shamim):

c=?

OpenStudy (anonymous):

y+2=1/2x + 1/2 -2 -2 y=1/2x -3/2 ?

OpenStudy (anonymous):

c= 3/2?

OpenStudy (anonymous):

-3/2 *

OpenStudy (shamim):

ok. u r right

OpenStudy (anonymous):

Thanks guys

OpenStudy (shamim):

now u hv to put the value of m nd c into y=mx+c

OpenStudy (shamim):

put m=1/2 c=-3/2

OpenStudy (shamim):

tell me ur result

OpenStudy (shamim):

u r most welcome